Whosyourdaddy:

how do you do x/1 over 5/(x-1)

XxXNessalulbaddieXxX:

u got like a screen shot

XxXNessalulbaddieXxX:

@whosyourdaddy wrote:
sry my arrows over the 1
yea lmao

Whosyourdaddy:

snowflake0531:

\[\frac{ x \div 1}{ \frac{ 5}{x-1 } } =\frac{ x }{ \frac{5}{x-1}}= x(\frac{x-1}{5}) \] From there, you can multiply the numerator, which you would get \(\frac{x^2-x}{5}\)

Debrus:

@snowflake0531 wrote:
\[\frac{ x \div 1}{ \frac{ 5}{x-1 } } =\frac{ x }{ \frac{5}{x-1}}= x(\frac{x-1}{5}) \]
note: \( \frac{ x \div 1}{ \frac{ 5}{x-1 } } =\frac{ x }{ \frac{5}{x-1}}= \frac{\color{red}{x(x-1)}}{5} \text{ and then youd get } \frac{x^2-x}{5} \)